# Low-energy electron scattering from aminoacetonitrile using the R-matrix method

### v1

@@ Abstract
Aminoacetonitrile occupies a prime importance in the interface between astrochemistry and prebiotic chemistry. Its detection in the ISM establishes it as part of the organic inventory of star-forming regions, while its role as a glycine precursor highlights its significance for origins-of-life scenarios. In this work, electron scattering from aminoacetonitrile has been studied using the $R$-matrix method in the low-energy range from 0 to 10 eV. The calculations were carried out within the $C_{s}$ point group using static-exchange (SE), static-exchange plus polarization (SEP), and configuration interaction (CI) models, with two basis sets (6-311G* and cc-pVTZ). Various scattering observables such as elastic, excitation, momentum transfer, and differential cross sections were examined. Since aminoacetonitrile is a prebiotically relevant molecule, these findings provide valuable insight into electron-driven processes in complex organic systems and form a theoretical foundation for future work on electron-induced reactivity in prebiotic and astrophysical environments.
@@
Prebiotic molecule , low-energy electron scattering , aminoacetonitrile , interstellar molecule , R-matrix method , elastic cross-section , momentum transfer cross-section , excitation cross section , differential cross section
@@
Low-energy electron scattering cross sections are reported for aminoacetonitrile.
R-matrix method is used, employing three different models (SE, SEP, and CI) and two different basis sets (6-311G* and cc-pVTZ).
Elastic, excitation, momentum transfer, and differential cross sections are computed over 0 to 10 eV.
Resonance features are observed, implying temporary negative ion formation.
@@ 1 Introduction
The importance of aminoacetonitrile lies in its prebiotic chemistry. In the classical Strecker synthesis, glycine is formed from ammonia, formaldehyde, and hydrogen cyanide via aminoacetonitrile as an intermediate (Koch et al., 2008). Theoretical studies show that this chemistry can occur on icy dust grains in cold interstellar environments, producing aminoacetonitrile under astrophysical conditions (Koch et al., 2008). Laboratory experiments confirm that both ultraviolet irradiation and simple thermal processing of ices containing ammonia and nitriles produce aminoacetonitrile (Danger et al., 2011b, a). Additional work has shown that aminoacetonitrile can form in water-rich ices and remains stable under moderate heating (Borget et al., 2012; Noble et al., 2013). Because glycine is difficult to detect directly in the interstellar medium (Snyder et al., 2005; Lattelais et al., 2011), aminoacetonitrile serves as a valuable alternative for understanding how far interstellar chemistry proceeds toward amino acid formation.
@@
Recent studies have highlighted the role of energetic processing in this chemistry. Low-energy electrons (LEEs, $\leq 20$ eV), created when cosmic rays ionize interstellar gas and ices, are especially important. They interact strongly with organic molecules, often through resonant processes that can lead either to fragmentation or to the production of reactive radicals (Boyer et al., 2016; Wu et al., 2023; Boudaıffa et al., 2000; Mason et al., 2014). In aminoacetonitrile, the presence of both an amino and a nitrile group introduces multiple reactive centers where LEEs may attach, potentially leading to selective bond cleavage, radical formation, or intramolecular rearrangements. Such processes could contribute both to the destruction of aminoacetonitrile in the ISM and to the generation of reactive intermediates that feed into further prebiotic chemistry. From a theoretical and experimental standpoint, studying electron scattering from aminoacetonitrile provides an opportunity to investigate how multifunctional prebiotic molecules respond to electron-driven processes. Unlike smaller nitriles (such as acetonitrile) or simple amines, aminoacetonitrile combines two chemically distinct functional groups, whose interplay may give rise to complex resonance structures and scattering behavior. Understanding these processes is critical for assessing the molecule’s stability under astrophysical conditions and for evaluating its prebiotic potential.
@@
For molecules similar to aminoacetonitrile, such as acetonitrile, low-energy electron experiments have revealed strong resonances and efficient dissociative electron attachment channels (Sailer et al., 2003; Maioli and Bettega, 2017; Zawadzki and Khakoo, 2018). Existing work on aminoacetonitrile focuses on dissociative electron attachment (DEA). Pelc et al. measured anion yields following low-energy electron collisions with aminoacetonitrile Pelc et al. (2016). These results show that LEEs can both destroy aminoacetonitrile and generate reactive fragments that may feed into further chemistry. Despite the progress, data on electron scattering and electron-induced processes from aminoacetonitrile are still limited. To date, no elastic or electronic excitation cross sections in the low energy range, i.e., below the ionization threshold, have been reported. Measurements of elastic and inelastic cross sections, resonance lifetimes, and fragmentation branching ratios are essential for refining astrochemical models. Without them, predictions of aminoacetonitrile abundances and lifetimes remain uncertain. Because the molecule contains both amino and nitrile groups, it is expected to support resonance channels that control its chemistry under electron impact. Thus, studying low-energy electron interactions with aminoacetonitrile is important not only for understanding its observed abundances but also for clarifying its role as a prebiotic precursor in interstellar, cometary, and planetary environments.
@@ 2 Theory
The UK R-matrix method (Tennyson, 2010), implemented through the Quantemol-N software (Tennyson et al., 2007), was used in this work to calculate low-energy electron scattering cross sections. The R-matrix approach is a well-established framework for studying electron–molecule collisions. It works by dividing configuration space into two regions separated by a spherical boundary of radius $a$, placed at the center of mass of the target molecule.
Inside the sphere, known as the inner region, the molecule and the incoming electron are treated together as an $(N+1)$-electrons system. In this region, short-range effects such as electron exchange and correlation dominate, so the system behaves like a bound state. Outside the sphere, i.e., in the outer region, the scattered electron no longer feels these short-range interactions. Instead, it interacts only with the long-range multipole potential of the target molecule. Here, the wavefunction of the scattering electron is expressed in a form that reduces the problem to a system of coupled differential equations, which can then be solved with standard numerical techniques.
The radius of the R-matrix sphere must be chosen carefully so that the short-range forces vanish at the boundary. In this study, a radius of $14a_{0}$ was used. The Hamiltonian of the $(N+1)$-electrons system is built separately for the inner and outer regions, and the solutions in these two regions are matched at the boundary of the sphere to ensure continuity of the wavefunctions.
@@
From the inner-region eigenfunctions, the R-matrix is constructed at the boundary. This R-matrix is then propagated outward and matched with asymptotic solutions obtained using the Gailitis expansion (Gailitis, 1976). Through this procedure, the $K$-matrix is determined. Using the POLYDCS code (Sanna and Gianturco, 1998), the $K$-matrix is converted into the $S$- and $T$-matrices, which are directly related to the scattering observables. The scattering cross sections are finally calculated from the elements of the $T$-matrix.
To examine the scattering under different levels of approximation, three models were used: the static exchange (SE), static exchange plus polarization (SEP), and configuration interaction (CI) models. The SE model provides a simple baseline, where the target molecule is represented by a frozen Hartree–Fock potential. The SEP model improves on this by including the distortion of the target’s electron cloud due to the incident electron, thus accounting for polarization effects. The CI model gives the most accurate description by also including excitations of the target, which are essential for capturing resonance effects and providing reliable excitation dynamics.
@@ 2.1 Present Target State
The point group for aminoacetonitrile is $C_{s}$. For the present calculation, the geometrical parameters for ground state aminoacetonitrile have been taken from the CCCBDB database (NIST, n.d.). The basis sets used to construct the target wavefunctions are 6-311G* and cc-pVTZ. The scattering calculations are done using static exchange (SE), static exchange plus polarization (SEP), and configuration interaction (CI) models. The ground state electronic configuration is given as:
@@ 3 Results and Discussions
This section depicts the results obtained for electron-aminoacetonitrile scattering in the energy range 0 to 10 eV. The results are presented in graphical formats.
@@ 3.1 Eigenphase sum
The eigenphase sums for electron scattering from aminoacetonitrile in the $C_{s}$ point group were obtained within the $R$-matrix framework for two irreducible representations, $^{2}A^{\prime}$ and $^{2}A^{\prime\prime}$. Calculations were performed with three target models—static-exchange (SE), static-exchange plus polarization (SEP), and configuration interaction (CI) using two different basis sets (6-311G* and cc-pVTZ), with an $R$-matrix radius of $12\,a_{0}$. The results are presented in figure 2.
In the static-exchange (SE) approximation, the scattering eigenphases are relatively smooth, with only weak evidence for sharp resonance behavior. For the 6-311G* basis, a single broad structure was identified in the $^{2}A^{\prime\prime}$ channel at 5.70 eV with a width of 1.98 eV. The corresponding eigenphase sum shows only a gradual rise, lacking the steep $\pi$-like jump characteristic of well-defined resonances. With the cc-pVTZ basis, a similar feature is found slightly higher in energy at 6.45 eV with a broader width of 2.71 eV. No clear resonant structure was detected in the $^{2}A^{\prime}$ symmetry for either basis set at the SE level. Overall, the SE model yields only diffuse, basis-dependent structures without strong eigenphase support for genuine resonances.
When target polarization is included in the static-exchange-plus-polarization (SEP) model, the situation changes noticeably. For 6-311G*, two resonances emerge: a $^{2}A^{\prime}$ state at 5.09 eV with a width of 1.80 eV and a $^{2}A^{\prime\prime}$ state at 4.75 eV with a narrower width of 1.20 eV. The eigenphase sums in this case show sharper, arctangent-like rises around these energies, confirming their resonant nature. Using the cc-pVTZ basis yields nearly identical results: the $^{2}A^{\prime}$ resonance appears at 5.15 eV with a width of 1.84 eV, while the $^{2}A^{\prime\prime}$ resonance is located at 4.81 eV with a width of 1.25 eV. Thus, polarization stabilizes and lowers the resonance energies compared to SE and produces consistent results across basis sets.
@@
Overall, the progression across models and basis sets is consistent. SE provides only broad, poorly defined features; SEP lowers and sharpens the first resonances into the 4–5 eV region with improved eigenphase support; and CI yields robust low-energy resonances at $\approx$4.3–4.6 eV in both symmetries, with widths of $\sim$1–1.7 eV. The larger cc-pVTZ basis generally shifts SE resonances upward and produces marginally narrower SEP and CI widths, but the low-energy CI resonances converge very well between the two basis sets. The additional high-energy, long-lived CI resonances at 7–10 eV appear in both basis sets, though with slight variations in energy and width. These features are highly sensitive and may correspond either to real narrow states or to artifacts of the fitting, requiring further validation through direct inspection of eigenphase sums and cross sections.
@@ 3.2 Elastic cross section
Figure 3 represents the elastic electron scattering cross sections of aminoacetonitrile, calculated within the $R$-matrix framework. The results show systematic dependence on both the chosen basis set and the treatment of electronic correlation. For both the 6-311G* and cc-pVTZ basis sets, the cross sections display the characteristic steep rise at low energies, consistent with the long-range polarization interaction, followed by a gradual decrease with increasing electron energy. This overall trend reflects the dominance of $s$-wave scattering near the threshold.
The most pronounced model dependence occurs in the 4–5 eV region, where the inclusion of polarization or correlation modifies both the shape and magnitude of the cross section. At the static-exchange (SE) level the elastic cross section is smooth. This behavior is consistent with the eigenphase sums, which showed only weak, broad features in the SE model without clear resonant jumps. In contrast, when target polarization is included in the SEP model, the cross sections exhibit a distinct resonant enhancement near 4.8–5.1 eV. This structure is observed for both basis sets, although it is slightly more pronounced in the cc-pVTZ calculation and corresponds directly to the low-lying resonances identified in the eigenphase analysis.
At the configuration-interaction (CI) level, the eigenphase analysis revealed robust resonances in both symmetries at 4.3–4.6 eV with widths of about 1–1.7 eV. In the elastic cross sections these resonances manifest as modest shoulders or plateaux, rather than sharp peaks. The overall magnitude of the CI cross sections in this energy region is reduced compared to SE and SEP, which suggests that correlation effects redistribute scattering flux in a way that sharpens the eigenphase behavior but does not strongly increase the total elastic signal. This is a common effect in polyatomic systems, where correlation improves the resonance description but also couples elastic and inelastic channels, thereby lowering the integrated elastic yield.
The comparison between basis sets shows that 6-311G* and cc-pVTZ give very similar elastic cross sections once correlation is included. At the SEP level, cc-pVTZ yields a slightly more pronounced resonant shoulder near 5 eV, while at the CI level, both basis sets produce nearly identical curves across the full energy range. This agreement indicates that the CI/cc-pVTZ model provides the most reliable description of the low-energy resonances and the corresponding elastic scattering behavior. The narrow high-energy resonances seen in the CI eigenphases between 7 and 10 eV do not contribute prominently to the total elastic cross section, consistent with their very small fitted widths and likely weak coupling. The convergence of the two basis sets at the CI level reinforces the robustness of these low-energy resonances, which are therefore assigned as genuine scattering states of aminoacetonitrile.

### v2

# Electron scattering from aminoacetonitrile: effects of polarisation-correlation and basis-set on cross section

@@ Abstract
Aminoacetonitrile occupies a prime importance in the interface between astrochemistry and prebiotic chemistry. Its detection in the ISM establishes it as part of the organic inventory of star-forming regions, while its role as a glycine precursor highlights its significance for origins-of-life scenarios. In this work, electron scattering from aminoacetonitrile has been studied using the $R$-matrix method in the low-energy range from $\sim$0 to 10 eV. The calculations were carried out within the $C_{s}$ point group using static-exchange (SE), static-exchange plus polarisation (SEP), and configuration interaction (CI) models, with two basis sets (6-311G* and cc-pVTZ) to understand their dependence on cross section. Various scattering observables, such as differential elastic cross section, integral elastic, excitation, and momentum transfer cross sections, were examined. Since aminoacetonitrile is a prebiotically relevant molecule, these findings provide valuable insight into electron-driven processes in complex organic systems and form a theoretical foundation for future work on electron-induced reactivity in prebiotic and astrophysical environments.
@@
Prebiotic molecule , low-energy electron scattering , aminoacetonitrile , R-matrix method , elastic cross-section , excitation cross section , differential cross section , polarisation-correlation
@@
Low-energy (below 10 eV) electron scattering from the aminoacetonitrile molecule is investigated here.
Sensitivity studies were conducted employing three models (SE, SEP, and CI) and two different basis sets (6-311G* and cc-pVTZ) under the R-matrix method.
Differential elastic and integral elastic, excitation, and momentum transfer cross sections are computed on a fine energy grid.
Resonance features are observed, implying temporary negative ion formation during collision.
@@ 1 Introduction
The importance of aminoacetonitrile lies in its prebiotic chemistry. In the classical Strecker synthesis, glycine is formed from ammonia, formaldehyde, and hydrogen cyanide via aminoacetonitrile as an intermediate (koch2008theoretical). Theoretical studies show that this chemistry can occur on icy dust grains in cold interstellar environments, producing aminoacetonitrile under astrophysical conditions (koch2008theoretical). Laboratory experiments confirm that both ultraviolet irradiation and simple thermal processing of ices containing ammonia and nitriles produce aminoacetonitrile (danger2011experimental; danger2011experimentalb). Additional work has shown that aminoacetonitrile can form in water-rich ices and remains stable under moderate heating (borget2012aminoacetonitrile; noble2013thermal). As glycine is difficult to detect directly in the interstellar medium (snyder2005rigorous; lattelais2011detectability), aminoacetonitrile serves as a valuable alternative for understanding how far interstellar chemistry proceeds toward amino acid formation.
@@
Recent studies have highlighted the role of energetic processing in this chemistry. Low-energy electrons (LEEs, $\leq 20$ eV), created when cosmic rays ionise interstellar gas and ices, are especially important. They interact strongly with organic molecules, often through resonant processes that can lead either to fragmentation or to the production of reactive radicals (boyer2016role; wu2023role; boudaiffa2000resonant; mason2014electron). In aminoacetonitrile, the presence of both an amino and a nitrile group introduces multiple reactive centres where LEEs may attach, potentially leading to selective bond cleavage, radical formation, or intramolecular rearrangements. Such processes could contribute both to the destruction of aminoacetonitrile in the ISM and to the generation of reactive intermediates that feed into further prebiotic chemistry. From a theoretical and experimental standpoint, studying electron scattering from aminoacetonitrile provides an opportunity to investigate how multifunctional prebiotic molecules respond to electron-driven processes. Unlike smaller nitriles (such as acetonitrile) or simple amines, aminoacetonitrile combines two chemically distinct functional groups, whose interplay may give rise to complex resonance structures and scattering behaviour. Understanding these processes is critical for assessing the molecule’s stability under astrophysical conditions and for evaluating its prebiotic potential.
@@
For molecules similar to aminoacetonitrile, such as acetonitrile, low-energy electron experiments have revealed strong resonances and efficient dissociative electron attachment channels (sailer2003low; maioli2017low; zawadzki2018low). Existing work on aminoacetonitrile focuses on dissociative electron attachment (DEA). Pelc et al. measured anion yields following low-energy electron collisions with aminoacetonitrile (pelc2016formation). These results show that LEEs can both destroy aminoacetonitrile and generate reactive fragments that may feed into further chemistry. Despite the progress, data on electron scattering and electron-induced processes from aminoacetonitrile are still limited. To date, no elastic or electronic excitation cross sections in the low energy range, i.e., below the ionisation threshold, have been reported. Measurements of elastic and inelastic cross sections, resonance lifetimes, and fragmentation branching ratios are essential for refining astrochemical models. Without them, predictions of aminoacetonitrile abundances and lifetimes remain uncertain. Because the molecule contains both amino and nitrile groups, it is expected to support resonance channels that control its chemistry under electron impact. Thus, studying low-energy electron interactions with aminoacetonitrile is important not only for understanding its observed abundances but also for clarifying its role as a prebiotic precursor in interstellar, cometary, and planetary environments.
@@ 2 Theory
The UK R-matrix method (tennyson2010electron), implemented through the Quantemol-N software (tennyson2007quantemol), was used in this work to calculate low-energy electron scattering cross sections. The R-matrix approach is a well-established framework for studying electron–molecule collisions. It works by dividing configuration space into two regions separated by a spherical boundary of radius $a$, placed at the center of mass of the target molecule.
Inside the sphere, known as the inner region, the molecule and the incoming electron are treated together as an $(N+1)$ electron system. In this region, short-range effects such as electron exchange and correlation dominate, so the system behaves like a bound state. Outside the sphere (outer region), the scattered electron no longer feels these short-range interactions. Instead, it interacts only with the long-range multipole potential of the target molecule. Here, the wavefunction of the scattering electron is expressed in a form that reduces the problem to a system of coupled differential equations, which can then be solved with standard numerical techniques.
The radius of the R-matrix sphere must be chosen carefully so that the short-range forces vanish at the boundary. In this study, a radius of $12a_{0}$ was used. The Hamiltonian of the $(N+1)$ electron system is built separately for the inner and outer regions, and the solutions in these two regions are matched at the boundary of the sphere to ensure continuity of the wavefunctions.
@@
From the inner-region eigenfunctions, the R-matrix is constructed at the boundary. This R-matrix is then propagated outward and matched with asymptotic solutions obtained using the Gailitis expansion (gailitis1976new). Through this procedure, the $K$-matrix is determined. Using the POLYDCS code (sanna1998differential), the $K$-matrix is converted into the $S$- and $T$-matrices, which are directly related to the scattering observables. The scattering cross sections are finally calculated from the elements of the $T$-matrix.
To examine the scattering under different levels of approximation, three models were used: the static exchange (SE), static exchange plus polarisation (SEP), and configuration interaction (CI) models. The SE model provides a simple baseline, where the target molecule is represented by a frozen Hartree–Fock potential. The SEP model improves on this by including the distortion of the target’s electron cloud due to the incident electron, thus accounting for polarisation effects. The CI model gives the most accurate description by also including excitations of the target, which are essential for capturing resonance effects and providing reliable excitation dynamics.
@@ 2.1 Present Target State
The point group for aminoacetonitrile is $C_{s}$. For the present calculation, the geometrical parameters for ground state aminoacetonitrile have been taken from the CCCBDB database (CCCBDB). The basis sets used to construct the target wavefunctions are 6-311G* and cc-pVTZ. The scattering calculations are done using static exchange (SE), static exchange plus polarisation (SEP), and configuration interaction (CI) models. The ground state electronic configuration is given as:
@@ 3 Results and Discussions
This section depicts the results obtained for electron-aminoacetonitrile scattering in the energy range $\sim$0 to 10 eV. The results are presented in graphical formats.
@@ 3.1 Eigenphase sum
The eigenphase sums for electron scattering from aminoacetonitrile in the $C_{s}$ point group were obtained within the $R$-matrix framework for two irreducible representations, $^{2}A^{\prime}$ and $^{2}A^{\prime\prime}$. Calculations were performed with three target models: static-exchange (SE), static-exchange plus polarisation (SEP), and configuration interaction (CI) using two different basis sets (6-311G* and cc-pVTZ), with an $R$-matrix radius of $12\,a_{0}$. The results are presented in figure 2.
In the static-exchange (SE) approximation, the scattering eigenphases are relatively smooth, with only weak evidence for sharp resonance behaviour. For the 6-311G* basis, a single broad structure was identified in the $^{2}A^{\prime\prime}$ channel at 5.70 eV with a width of 1.98 eV. The corresponding eigenphase sum shows only a gradual rise, lacking the steep $\pi$-like jump characteristic of well-defined resonances. With the cc-pVTZ basis, a similar feature is found slightly higher in energy at 6.45 eV with a broader width of 2.71 eV. No clear resonant structure was detected in the $^{2}A^{\prime}$ symmetry for either basis set at the SE level. Overall, the SE model yields only diffuse, basis-dependent structures without strong eigenphase support for genuine resonances.
When target polarisation is included in the static-exchange-plus-polarisation (SEP) model, the situation changes noticeably. For 6-311G*, two resonances emerge: a $^{2}A^{\prime}$ state at 5.09 eV with a width of 1.80 eV and a $^{2}A^{\prime\prime}$ state at 4.75 eV with a narrower width of 1.20 eV. The eigenphase sums in this case show sharper, arctangent-like rises around these energies, confirming their resonant nature. Using the cc-pVTZ basis yields nearly identical results: the $^{2}A^{\prime}$ resonance appears at 5.15 eV with a width of 1.84 eV, while the $^{2}A^{\prime\prime}$ resonance is located at 4.81 eV with a width of 1.25 eV. Thus, polarisation stabilises and lowers the resonance energies compared to SE and produces consistent results across basis sets.
@@
Overall, the progression across models and basis sets is consistent. SE provides only broad, poorly defined features; SEP lowers and sharpens the first resonances into the 4–5 eV region with improved eigenphase support; and CI yields robust low-energy resonances at $\sim$4.3–4.6 eV in both symmetries, with widths of $\sim$1–1.7 eV. The larger cc-pVTZ basis generally shifts SE resonances upward and produces marginally narrower SEP and CI widths, but the low-energy CI resonances converge very well between the two basis sets. The additional high-energy, long-lived CI resonances at 7–10 eV appear in both basis sets, though with slight variations in energy and width. These features are highly sensitive and may correspond either to real narrow states or to artifacts of the fitting, requiring further validation through direct inspection of eigenphase sums and cross sections.
@@ 3.2 Elastic cross section
Figure 3 represents the integral elastic electron scattering cross sections of aminoacetonitrile for various models and basis sets. The results show systematic dependence on both the chosen basis set and the treatment of electronic correlation. For both the 6-311G* and cc-pVTZ basis sets, the cross sections display the characteristic steep rise at low energies, consistent with the long-range polarisation interaction, followed by a gradual decrease with increasing electron energy. This overall trend reflects the dominance of $s$-wave scattering near the threshold.
The most pronounced model dependence occurs in the 4–5 eV region, where the inclusion of polarisation and correlation modifies both the shape and magnitude of the cross section. At the static-exchange (SE) level the elastic cross section is smooth. This behaviour is consistent with the eigenphase sums, which showed only weak, broad features in the SE model without clear resonant jumps. In contrast, when target polarisation is included in the SEP model, the cross sections exhibit a distinct resonant enhancement near 4.8–5.1 eV. This structure is observed for both basis sets, although it is slightly more pronounced in the cc-pVTZ calculation and corresponds directly to the low-lying resonances identified in the eigenphase analysis.
At the configuration-interaction (CI) level, the eigenphase analysis revealed robust resonances in both symmetries at 4.3–4.6 eV with widths of about 1–1.7 eV. In the elastic cross sections these resonances manifest as modest shoulders or plateaux, rather than sharp peaks. The overall magnitude of the CI cross sections in this energy region is reduced compared to SE and SEP, which suggests that correlation effects redistribute scattering flux in a way that sharpens the eigenphase behaviour but does not strongly increase the total elastic signal. This is a common effect in polyatomic systems, where correlation improves the resonance description but also couples elastic and inelastic channels, thereby lowering the integrated elastic yield.
The comparison between basis sets shows that 6-311G* and cc-pVTZ give very similar elastic cross sections once correlation is included. At the SEP level, cc-pVTZ yields a slightly more pronounced resonant shoulder near 5 eV, while at the CI level, both basis sets produce nearly identical curves across the full energy range. This agreement indicates that the CI/cc-pVTZ model provides the most reliable description of the low-energy resonances and the corresponding elastic scattering behaviour. The narrow high-energy resonances seen in the CI eigenphases between 7 and 10 eV do not contribute prominently to the total elastic cross section, consistent with their very small fitted widths and likely weak coupling. The convergence of the two basis sets at the CI level reinforces the robustness of these low-energy resonances, which are therefore assigned as genuine scattering states of aminoacetonitrile.